# The Addition Theorem for the Algebraic Entropy of Torsion Nilpotent Groups

### v2

@@ Abstract.
The Addition Theorem for the algebraic entropy of group endomorphisms of torsion abelian groups was proved by Dikranjan, Goldsmith, Salce and Zanardo [4]. It was later extended by Shlossberg [10] to torsion nilpotent groups of class $2$. As our main result, we prove the Addition Theorem for endomorphisms of torsion nilpotent groups of arbitrary nilpotency class. As an application, we show that if $G$ is a torsion nilpotent group, then for every $\phi\in\operatorname{End}(G)$ either the entropy $h(\phi)$ is infinite or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}}$. We further obtain, for automorphisms of locally finite groups, the Addition Theorem with respect to all terms of the upper central series; in particular, the Addition Theorem holds for automorphisms of $\omega$-hypercentral groups. Finally, we establish a reduction principle: if $\mathfrak{X}$ is a variety of locally finite groups, then the Addition Theorem for endomorphisms holds in $\mathfrak{X}$ if and only if it holds for locally finite groups generated by bounded sets.
@@ 1. Introduction
As a first application, we obtain a dichotomy for the algebraic entropy of endomorphisms of torsion nilpotent groups: if $G$ is torsion nilpotent and $\phi\in\operatorname{End}(G)$ has finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}}$.
We then turn to locally finite groups. Using the nilpotent case as a key input, we show that for every locally finite group $G$ and every automorphism $\varphi\in\operatorname{Aut}(G)$, the Addition Theorem holds with respect to each term $Z_{n}(G)$ of the upper central series. As a consequence, in Section 5 we show that the Addition Theorem holds for automorphisms of $\omega$-hypercentral groups.
Finally, we establish in Section 5 a reduction principle for varieties of locally finite groups, showing that the Addition Theorem for endomorphisms holds in such a variety if and only if it holds for locally finite groups generated by bounded sets.
@@ 2.1. Notation and terminology
A group $G$ is called solvable if there exists $k\in{\mathbb{N}}_{+}$ such that $G^{(k)}=\{1\}$. In particular, $G$ is metabelian if $G^{\prime\prime}=G^{(2)}=\{1\}$. We also write $\gamma_{1}(G)=G$, $\gamma_{i+1}(G)=[\gamma_{i}(G),G]$ for the lower central series. In particular, $G^{\prime}=\gamma_{2}(G)$ and if $G$ is nilpotent of class at most $n$, then $\gamma_{n+1}(G)=\{1\}$ and $\gamma_{n}(G)\leq Z(G)$, where $Z(G)$ denotes the center of $G$.
@@
If $\phi\in\operatorname{End}(G)$, then a subgroup $H$ of $G$ is called $\phi$-invariant if $\phi(H)\subseteq H$, and $H$ is $\phi$-stable if $\phi\in\operatorname{Aut}(G)$ and $\phi(H)=H$. A subgroup $H$ of $G$ is characteristic if $H$ is $\phi$-invariant for every $\phi\in\operatorname{Aut}(G)$, and fully characteristic if the same holds for every $\phi\in\operatorname{End}(G)$. For every $n$ we let $G[n]$ be the fully characteristic subgroup of $G$ generated by the set $\{x\in G:x^{n}=e\}$. We say that a subset $F\subseteq G$ is bounded if $F\subseteq G[n]$ for some $n\in\mathbb{N}$, equivalently, if the orders of the elements of $F$ are uniformly bounded.
@@ 3. Auxiliary results
Let $\mathfrak{X}$ be a variety of torsion solvable groups. If the Addition Theorem holds in $\mathfrak{X}$ for endomorphisms, then for every $G\in\mathfrak{X}$ and every $\phi\in\operatorname{End}(G)$ either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}}.$
@@
where $\bar{\phi}\in\operatorname{End}(G/G^{\prime})$ is the induced endomorphism. Since $G^{\prime}$ is abelian and $G/G^{\prime}$ is solvable of class less than $n$, the induction hypothesis implies that there exist $\alpha,\beta\in{\mathbb{N}}$ such that $h(\phi\upharpoonright_{G^{\prime}})=\log(\alpha)$ and $h(\bar{\phi})=\log(\beta)$. Therefore
@@ 4. Proving the main result
If $G$ is a torsion nilpotent group, and $\phi\in\operatorname{End}(G)$ is of finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}}.$
@@
Proposition 4.8 (Addition Theorem for the $n$-center for automorphisms).
Let $G$ be a locally finite group and let $\varphi\in\operatorname{Aut}(G)$. Then for every $n\in\mathbb{N}$,
@@
In other words, $\operatorname{AT}(G,\varphi,Z_{n}(G))$ holds for every $n\in\mathbb{N}$.
@@ 5. A reduction principle for locally finite groups
Let $G$ be a locally finite group. If $\operatorname{AT}(G[n!])$ holds for endomorphisms for every $n\in\mathbb{N}$, then $\operatorname{AT}(G)$ holds for endomorphisms.
@@
For a locally finite group $G$, the family $\{G[n!]\}_{n\in\mathbb{N}}$ forms a chain of fully characteristic subgroups whose union is $G$, so the conclusion follows from Proposition 5.21. ∎
@@
Moreover, either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}}.$
@@
Since $G$ is $\omega$-hypercentral, we have $G=\bigcup_{n\in\mathbb{N}}Z_{n}(G)$. Each $Z_{n}(G)$ is a characteristic (hence $\phi$-stable) torsion nilpotent subgroup of $G$. By Theorem 4.5, $\operatorname{AT}(Z_{n}(G))$ holds for endomorphisms, and hence also for automorphisms. Therefore Proposition 5.22 yields $\operatorname{AT}(G)$ for automorphisms.
For the entropy dichotomy, note that a torsion $\omega$-hypercentral group is locally nilpotent and hence locally finite. Applying Proposition 5.1 to the directed chain $\{Z_{n}(G)\}_{n\in\mathbb{N}}$ we obtain
$h(\phi)=\sup_{n\in\mathbb{N}}h\bigl(\phi\!\upharpoonright_{Z_{n}(G)}\bigr).$
By Corollary 4.7, for each $n$ either $h(\phi\!\upharpoonright_{Z_{n}(G)})=\infty$ or $h(\phi\!\upharpoonright_{Z_{n}(G)})=\log(\alpha_{n})$ for some $\alpha_{n}\in{\mathbb{N}}$. If $h(\phi)<\infty$, then all these entropies are finite, so
$h(\phi)=\sup_{n\in\mathbb{N}}\log(\alpha_{n})=\log(\alpha),$
where $\alpha=\sup_{n\in\mathbb{N}}\alpha_{n}\in{\mathbb{N}}$. ∎

### v3

@@ Abstract.
The Addition Theorem for the algebraic entropy of group endomorphisms of torsion abelian groups was proved by Dikranjan, Goldsmith, Salce and Zanardo [4]. It was later extended by Shlossberg [10] to torsion nilpotent groups of class $2$. As our main result, we prove the Addition Theorem for endomorphisms of torsion nilpotent groups of arbitrary nilpotency class. As an application, we show that if $G$ is a torsion nilpotent group, then for every $\phi\in\operatorname{End}(G)$ either the entropy is infinite or $h(\phi)=\log(\alpha)$ for a positive integer $\alpha$. We further obtain, for automorphisms of locally finite groups, the Addition Theorem with respect to every term of the upper central series; in particular, it holds for automorphisms of $\omega$-hypercentral groups. Finally, we establish a reduction principle: if $\mathfrak{X}$ is a class of locally finite groups closed under taking subgroups and quotients, then the Addition Theorem for endomorphisms holds in $\mathfrak{X}$ if and only if it holds for locally finite groups generated by bounded sets.
@@ 1. Introduction
As a first application, we obtain a dichotomy for the algebraic entropy of endomorphisms of torsion nilpotent groups: if $G$ is torsion nilpotent and $\phi\in\operatorname{End}(G)$ has finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}$.
We then turn to general locally finite groups. Using the nilpotent case as a key input, we show that for every locally finite group $G$ and every automorphism $\varphi\in\operatorname{Aut}(G)$, the Addition Theorem holds with respect to each term $Z_{n}(G)$ of the upper central series. As a consequence, in Section 5 we show that the Addition Theorem holds for automorphisms of $\omega$-hypercentral groups.
Finally, we establish in Section 5 a reduction principle for classes of locally finite groups closed under taking subgroups and quotients, showing that the Addition Theorem for endomorphisms holds in such a class if and only if it holds for locally finite groups generated by bounded sets.
@@ 2.1. Notation and terminology
A group $G$ is called solvable if there exists $k\in{\mathbb{N}}$ such that $G^{(k)}=\{1\}$. In particular, $G$ is metabelian if $G^{\prime\prime}=G^{(2)}=\{1\}$. We also write $\gamma_{1}(G)=G$, $\gamma_{i+1}(G)=[\gamma_{i}(G),G]$ for the lower central series. In particular, $G^{\prime}=\gamma_{2}(G)$ and if $G$ is nilpotent of class at most $n$, then $\gamma_{n+1}(G)=\{1\}$ and $\gamma_{n}(G)\leq Z(G)$, where $Z(G)$ denotes the center of $G$.
@@
If $\phi\in\operatorname{End}(G)$, then a subgroup $H$ of $G$ is called $\phi$-invariant if $\phi(H)\subseteq H$, and $H$ is $\phi$-stable if $\phi\in\operatorname{Aut}(G)$ and $\phi(H)=H$. A subgroup $H$ of $G$ is characteristic if $H$ is $\phi$-invariant for every $\phi\in\operatorname{Aut}(G)$, and fully characteristic if the same holds for every $\phi\in\operatorname{End}(G)$. For every $n$ we let $G[n]$ be the fully characteristic subgroup of $G$ generated by the set $\{x\in G:x^{n}=e\}$. We say that a subset $F\subseteq G$ is bounded if the orders of the elements of $F$ are uniformly bounded.
@@ 3. Auxiliary results
Let $\mathfrak{X}$ be a class of torsion solvable groups closed under taking subgroups and quotients. If the Addition Theorem holds in $\mathfrak{X}$ for endomorphisms, then for every $G\in\mathfrak{X}$ and every $\phi\in\operatorname{End}(G)$ either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
@@
where $\bar{\phi}\in\operatorname{End}(G/G^{\prime})$ is the induced endomorphism. Since $G^{\prime}$ is abelian and $G/G^{\prime}$ is solvable of class less than $n$, the induction hypothesis implies that there exist $\alpha,\beta\in{\mathbb{N}_{+}}$ such that $h(\phi\upharpoonright_{G^{\prime}})=\log(\alpha)$ and $h(\bar{\phi})=\log(\beta)$. Therefore
@@ 4. Proving the main result
If $G$ is a torsion nilpotent group, and $\phi\in\operatorname{End}(G)$ is of finite entropy, then $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
@@
Proposition 4.8 (Addition Theorem for the $n$-center of automorphisms).
Let $G$ be a locally finite group and let $\varphi\in\operatorname{Aut}(G)$. Then for every $n\in{\mathbb{N}_{+}}$,
@@
In other words, $\operatorname{AT}(G,\varphi,Z_{n}(G))$ holds for every $n\in{\mathbb{N}_{+}}$.
@@ 5. A reduction principle for locally finite groups
Let $G$ be a locally finite group. If $\operatorname{AT}(G[n!])$ holds for endomorphisms for every $n\in{\mathbb{N}_{+}}$, then $\operatorname{AT}(G)$ holds for endomorphisms.
@@
For a locally finite group $G$, the family $\{G[n!]\}_{n\in{\mathbb{N}_{+}}}$ forms a chain of fully characteristic subgroups whose union is $G$, so the conclusion follows from Proposition 5.21. ∎
@@
Moreover, either $h(\phi)=\infty$ or $h(\phi)=\log(\alpha)$ for some $\alpha\in{\mathbb{N}_{+}}.$
@@
Since $G$ is $\omega$-hypercentral, we have $G=\bigcup_{n\in{\mathbb{N}_{+}}}Z_{n}(G)$. Each $Z_{n}(G)$ is a characteristic (hence $\phi$-stable) torsion nilpotent subgroup of $G$. By Theorem 4.5, $\operatorname{AT}(Z_{n}(G))$ holds for endomorphisms, and hence also for automorphisms. Therefore Proposition 5.22 yields $\operatorname{AT}(G)$ for automorphisms.
For the entropy dichotomy, note that a torsion $\omega$-hypercentral group is locally nilpotent and hence locally finite. Applying Proposition 5.1 to the directed chain $\{Z_{n}(G)\}_{n\in{\mathbb{N}_{+}}}$ we obtain
$h(\phi)=\sup_{n\in{\mathbb{N}_{+}}}h\bigl(\phi\!\upharpoonright_{Z_{n}(G)}\bigr).$
By Corollary 4.7, for each $n$ either $h(\phi\!\upharpoonright_{Z_{n}(G)})=\infty$ or $h(\phi\!\upharpoonright_{Z_{n}(G)})=\log(\alpha_{n})$ for some $\alpha_{n}\in{\mathbb{N}_{+}}$. If $h(\phi)<\infty$, then all these entropies are finite, so
$h(\phi)=\sup_{n\in{\mathbb{N}_{+}}}\log(\alpha_{n})=\log(\alpha),$
where $\alpha=\sup_{n\in{\mathbb{N}_{+}}}\alpha_{n}\in{\mathbb{N}_{+}}$. ∎